# 3DSRBench: A Comprehensive 3D Spatial Reasoning Benchmark

### v1

@@ Abstract
3D spatial reasoning is the ability to analyze and interpret the positions, orientations, and spatial relationships of objects within the 3D space. This allows models to develop a comprehensive understanding of the 3D scene, enabling their applicability to a broader range of areas, such as autonomous navigation, robotics, and AR/VR. While large multi-modal models (LMMs) have achieved remarkable progress in a wide range of image and video understanding tasks, their capabilities to perform 3D spatial reasoning on diverse natural images are less studied. In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, with 2,772 manually annotated visual question-answer pairs across 12 question types. We conduct robust and thorough evaluation of 3D spatial reasoning capabilities by balancing the data distribution and adopting a novel FlipEval strategy. To further study the robustness of 3D spatial reasoning w.r.t. camera 3D viewpoints, our 3DSRBench includes two subsets with 3D spatial reasoning questions on paired images with common and uncommon viewpoints. We benchmark a wide range of open-sourced and proprietary LMMs, uncovering their limitations in various aspects of 3D awareness, such as height, orientation, location, and multi-object reasoning, as well as their degraded performance on images with uncommon camera viewpoints. Our 3DSRBench provide valuable findings and insights about the future development of LMMs with strong 3D reasoning capabilities. Our project page and dataset is available here.
@@ 1 Introduction
Recent large multi-modal models (LMMs) [1, 43, 3] have achieved significant improvements in a wide range of image and video understanding tasks, such as image captioning [2, 28], visual question answering [24, 20, 31, 10, 44], visual grounding [49], decision-making [34, 26, 7], and action recognition [48, 35]. However, recent studies have shown that even state-of-the-art LMMs exhibited limited 3D awareness [17, 36] and understanding of spatial relationships [46, 47], which are crucial for LMMs to develop a comprehensive grasp of the 3D scene, to enable high-level reasoning and interaction, and eventually, to achieve human-level intelligence.
Figure 1: 3D spatial reasoning performance comparison across leading models and human evaluation on 3DSRBench. Leading models (GPT-4o, Gemini-Pro, LLaVA-NeXT and Cambrian) lag behind human performance by at least -56% on average across four types of spatial relations (height, location, orientation, and multi-object). All models struggle with questions involving 3D orientation (e.g., “on the left”) and complex spatial reasoning (e.g., “viewpoint towards object”). See Fig. 2(a) for examples.
@@
Figure 2: Overview of the our 3DSRBench. (a) Example questions from the four main types of 3D spatial reasoning questions, i.e., height, location, orientation, and multi-object reasoning. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy to remove left/right biases in 3D with paired VQAs (see Sec. 3.4).
@@
To study the 3D spatial reasoning capabilities of LMMs, previous works exploited synthetic environments and generated images with 3D ground-truths [46, 47]. Visual question-answer pairs were automatically synthesized by applying pre-defined rules on the known 3D scene graphs and other object attributes. The synthetic images exhibit a significant domain gap with natural images and lacked the diversity and richness of contents in real-world. More recent works [13] explored real datasets with 3D annotations, e.g., Omni3D [6]. However, images in these datasets are still limited to specific domains, such as indoor rooms and self-driving scenes. In general, visual question-answer pairs generated with rule-based methods from 3D annotations (i) limit the scope of theirs datasets to a very small set of rigid objects, and (ii) cannot enable a fine-grained and robust evaluation of 3D spatial relationships that can only be achieved with human annotated datasets (see Sec. 3.1).
@@
In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, that features a wide variety of 3D spatial reasoning questions on diverse and open-vocabulary entities, including rigid objects, humans, animals, and implicit concepts, such as logo on a car or arrow on a billboard. We manually annotate 2,100 visual question-answer pairs on natural images from the MS-COCO dataset [29], covering 12 subtypes of questions from 4 main categories, i.e., height, location, orientation, and multi-object reasoning. Each category of questions focus on different combinations of 3D properties, such as object 3D location, 3D ground plane, camera extrinsic calibration, and/or object 3D poses. Examples from each question category are presented in Figure 2(a).
@@
Another challenge of 3D spatial reasoning arises from the 6D viewpoint of the camera, i.e., the 3D location and 3D orientation from which we are viewing the 3D scene. As shown in Fig. 4, 3D spatial reasoning questions can be easier for “common” 6D viewpoints, e.g., ones positioned at the eye level with natural viewing angles, while being more challenging for other “uncommon” viewpoints. Although uncommon viewpoints are less populated in most image datasets, cameras in embodied AI and robotics are often positioned in these uncommon viewpoints. Hence it is of crucial importance for LMMs to retain good 3D spatial reasoning performance for both common and uncommon viewpoints. To fairly compare the 3D spatial reasoning capabilities of LMMs w.r.t. different camera viewpoints, we annotate another 672 visual question-answer pairs on multi-view synthetic images rendered from the HSSD dataset [25].
@@
Besides benchmarking a wide variety of open-sourced and proprietary LMMs, our 3DSRBench serves as an important diagnosis benchmark for developing 3D spatially intelligent LMMs. Inspired by previous studies on 3D awareness of visual foundation models [17, 36], our 3DSRBench takes one step further and evaluates LMMs on fundamental 3D spatial reasoning questions, which provide valuable insights regarding the 3D awareness of visual encoders [39, 11, 21, 38, 27] and the 3D reasoning capabilities of language models [14, 50, 45, 16]. Such results would shed light on downstream tasks that build on 3D spatial reasoning, such as automatic navigation and robotic manipulation.
@@
To enable a comprehensive and robust evaluation of 3D spatial reasoning capabilities, our 3DSRBench adopts several key designs: (1) a balanced data distribution in multiple aspects, such as a roughly same number of yes/no answers and complementary images pairs that lead to opposite answers given the same question (see Fig. 2(b)); (2) avoiding questions with trivial answers; and (3) a novel FlipEval strategy designed for robust evaluation of 3D spatial reasoning questions.
@@
In summary, our 3DSRBench significantly advances the evaluation of 3D spatial reasoning capabilities by manually annotating 2,100 visual question-answer pairs on natural images from MS-COCO [29] and 672 on multi-view synthetic images. We annotate diverse and open-vocabulary entities and adopt key designs to enable a robust and thorough evaluation that cannot be done with existing datasets. We benchmark a wide variety of open-sourced and proprietary LMMs and study their 3D spatial reasoning capabilities w.r.t. different types of 3D awareness. We further analyze the robustness of 3D spatial reasoning capabilities w.r.t. common and uncommon camera 6D viewpoints and study a range of LMMs with different visual encoder designs. Experimental results on different splits of our 3DSRBench provide valuable findings and insights that will benefit future research on 3D spatially intelligent LMMs.
@@ 2 Related Works
3D spatial reasoning.
Previous studies on 3D spatial reasoning capabilities of vision-language models often adopt synthetic environments, e.g., Blender, with controllable simulation and 3D groundtruths for automatic question-answer generation [12, 46, 47]. However, synthetic images in these datasets exhibit a large domain gap with natural images and it remains unclear if insights and findings from these datasets would generalize to the real image domain. More recent works, such as SpatialRGPT [13] and Cambrian-1 [44], built on existing datasets with 3D annotations [6, 42, 5, 19, 41, 8] and generated visual question-answer pairs with pre-defined rules. Despite the improved image qualities, they are essentially limited to a small number of rigid object categories in Omni3D [6]. To enable a comprehensive and valuable evaluation of the 3D spatial reasoning capabilities, we manually annotate visual question-answer pairs, with diverse and open-vocabulary entities, such as logos on a car or arrows on the billboard, and questions about a comprehensive list of 3D spatial relationships from four main categories, i.e., height, location, orientation, and multi-object reasoning.
@@
In this section we introduce the 3DSRBench for comprehensively analyzing the 3D spatial reasoning capabilities of large multi-modal models (LMMs). We start by introducing the design considerations of 3DSRBench in Sec. 3.1, specifically how these design choices lead to a robust and valuable evaluation of 3D spatial reasoning capabilities. Then we present the four main question types in Sec. 3.2, as well as the challenges in each type of questions. Next we discuss the three splits of our 3DSRBench and their scopes in Sec. 3.3. In Sec. 3.4 we present our evaluation strategies, including CircularEval and FlipEval. Lastly in Sec. 3.5 we provide an overview of our data collection procedures and the summary statistics.
@@ 3.1 Design of 3DSRBench
When developing 3DSRBench, we incorporate the following four key designs to enable a robust and valuable evaluation of 3D spatial reasoning capabilities. First, our 3D spatial reasoning questions are based on open-vocabulary entities. Previous spatial reasoning benchmarks [9, 44] largely relied on existing datasets with 3D annotations [6], which limited their scope to a small number of rigid object categories. In our 3DSRBench, we annotate 3D spatial reasoning questions across a broad range of open-vocabulary entities (see Fig. 2), enabling a thorough analysis of the 3D awareness and 3D reasoning capabilities of LMMs over diverse, commonly encountered real-world objects. Next, we avoid questions with trivial answers. For instance, when comparing which of the two objects has a smaller 3D distance to a third anchor object, we avoid the cases when there is a significant gap between the two distances. We focus on samples where the two distances are close, but the correct answer can be derived with annotators’ consensus (see Sec. 3.5). Moreover, we implement a balanced data distribution in various aspects, such as a roughly same number of yes/no answers and complementary image pairs [20] that lead to opposite answers given the same 3D spatial reasoning question (see Fig. 2(b)). This effectively removes priors in the answer distribution, e.g., pedestrians are often located lower than street lights, or the fact that objects higher in 3D space are also higher in 2D image plane. This design ensures that models cannot exploit biases or shortcuts for a higher benchmark performance. Lastly, we adopt special evaluation strategies for robust evaluation, including previous CircularEval [33] and our novel FlipEval (see Sec. 3.4).
@@ 3.2 Question Types
(a) Height questions with different camera pitch rotations.
(b) Comparison between 2D and 3D spatial reasoning questions.
@@
Multi-object reasoning questions consider the 3D spatial relationships between multiple objects, such as asking which side of an object is facing another object, or with three objects, asking which of the given objects is facing towards or closer to the third object. In general, these questions require more advanced 3D awareness than simpler 3D concepts such as “closer” (to the camera) or “higher”, and require more complex reasoning over multiple objects’ 3D locations and orientations.
@@
Table 1: Overview of the 4 types of 3D spatial reasoning questions. Each type of 3D spatial reasoning questions require different combinations of 3D awareness, i.e., camera calibration, 3D location, 3D orientation, and multi-object reasoning.
@@ 3.4 Evaluation
3.5 Data Collection
We employ three annotators to annotate a total of 2,772 unique visual question-answer pairs across 12 question types. We follow the annotation principles as discussed in Sec. 3.1 and adopt an two-stage pipeline to ensure various criteria are met. Specifically after annotations are collected in the first stage, we review the quality of the collected data and reject samples with low quality or ones that lead to imbalanced data distribution. Further new annotations are collected if necessary. Furthermore, we collect human responses for all visual question-answer pairs and disregard samples that don’t reach consensus by human annotators.
@@ 4 Experiments
In this section, we start by introducing our experimental settings in Sec. 4.1. Then we benchmark the 3D spatial reasoning capabilities of various open-sourced and proprietary LMMs on our 3DSRBench in Sec. 4.2. Next we consider variants of LMMs with different visual encoders and study how the 3D awareness of visual encoders contribute to the final 3D spatial reasoning capabilities in Sec. 4.3. In Sec. 4.4 we evaluate various LMMs on our 3DSRBench-synthetic and analyze the robustness of LMMs w.r.t. uncommon camera viewpoints. Lastly we present some failure cases of GPT-4o in Sec. 4.5, revealing some interesting findings about the 3D spatial reasoning of LMMs.
@@ 4.1 Experimental Settings
We develop rule-based methods to augment the unique annotated visual question-answer pairs and obtain a larger number of testing data with a balanced and rich set of 3D spatial relationships. For instance, given a question asking which object is higher in the 3D world space, we generate a new question asking which object is lower in the 3D world space. We end up with 2,625 questions on MS-COCO images [29], i.e., 3DSRBench-real, and 846 questions on synthetic images, i.e., 3DSRBench-synthetic.
4.2 Results on 3D Spatial Reasoning Capabilities

### v2

@@ Abstract
3D spatial reasoning is the ability to analyze and interpret the positions, orientations, and spatial relationships of objects within the 3D space. This allows models to develop a comprehensive understanding of the 3D scene, enabling their applicability to a broader range of areas, such as autonomous navigation, robotics, and AR/VR. While large multi-modal models (LMMs) have achieved remarkable progress in a wide range of image and video understanding tasks, their capabilities to perform 3D spatial reasoning on diverse natural images are less studied. In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, with 2,772 manually annotated visual question-answer pairs across 12 question types. We conduct robust and thorough evaluation of 3D spatial reasoning capabilities by balancing the data distribution and adopting a novel FlipEval strategy. To further study the robustness of 3D spatial reasoning w.r.t. camera 3D viewpoints, our 3DSRBench includes two subsets with 3D spatial reasoning questions on paired images with common and uncommon viewpoints. We benchmark a wide range of open-sourced and proprietary LMMs, uncovering their limitations in various aspects of 3D awareness, such as height, orientation, location, and multi-object reasoning, as well as their degraded performance on images with uncommon camera viewpoints. Our 3DSRBench provide valuable findings and insights about future development of LMMs with strong 3D spatial reasoning abilities. Our project page and dataset is available here.
@@ 1 Introduction
Recent large multi-modal models (LMMs) [1, 48, 4] have achieved significant improvements in a wide range of image and video understanding tasks, such as image captioning [2, 35], visual question answering [28, 24, 39, 13, 52], visual grounding [58], decision making [42, 33, 10], and action recognition [57, 43]. Notably, the spatial reasoning ability [30, 28, 17, 54], i.e., parsing 2D and 3D spatial relationships between objects, serves as a crucial foundation for various high-level reasoning and interaction in downstream tasks. Studying the spatial reasoning ability of current LMMs will help us identify specific types of factual errors, uncover their fundamental limitations, and inform targeted improvements to further advance current LMMs.
Prior datasets [30, 28, 36, 31] studying spatial relationships often focused on relationships w.r.t. the viewer, e.g., object A is to the left of object B from the viewer’s perspective. We regard these as 2D spatial relationships as they can be captured merely from 2D bounding boxes of the objects (see Fig. 2(b)). They neglect 3D spatial relationships in the 3D world space or those from an object’s perspective. Capturing 3D spatial relationships between objects in the images would help LMMs understand and predict the interactions between objects, and enable a broader range of applications in 3D, e.g., robotics and embodied AI.
@@
Figure 1: Overview of our 3DSRBench. (a) Example questions from the four main types of 3D spatial reasoning questions, i.e., height, location, orientation, and multi-object reasoning. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy to remove left/right biases in 3D with paired VQAs (see Sec. 3.4).
@@
To study how LMMs can capture 3D spatial relationships, previous works often exploited synthetic environments and generated images with 3D ground-truths [55, 56]. Visual question-answer pairs were automatically synthesized by applying pre-defined rules to the known 3D scene graphs and object attributes. The synthetic images exhibit a significant domain gap with natural images and lacked the diversity and richness in real-world. More recent works [17] explored real datasets with 3D annotations, e.g., Omni3D [9]. However, images in these datasets are limited to specific domains, such as indoor rooms and self-driving scenes. In general, visual question-answer pairs generated with rule-based methods from 3D annotations (i) limit the scope of theirs datasets to a small set of rigid object categories, and (ii) cannot enable a fine-grained and robust evaluation of 3D spatial relationships that can only be achieved with human annotated datasets (see Sec. 3.1).
@@
In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, that features 2,772 3D spatial reasoning questions from 12 question types on diverse and open-vocabulary entities, including rigid objects, humans, animals, and implicit concepts, such as logo on a car or arrow on a billboard. We manually annotate 2,100 visual question-answer pairs on natural images from the MS-COCO dataset [37], covering 12 subtypes of questions from 4 main categories, i.e., height, location, orientation, and multi-object reasoning. Each category of questions focus on different combinations of 3D properties, such as object 3D location, 3D ground plane, camera extrinsic calibration, and/or object 3D poses. Examples from each question category are presented in Figure 1(a).
@@
Another challenge of 3D spatial reasoning arises from the 6D viewpoint of the camera, i.e., the 3D location and 3D orientation from which we are viewing the 3D scene. As shown in Fig. 3, 3D spatial reasoning questions can be easier for common 6D viewpoints, e.g., ones positioned at the eye level with natural viewing angles, while being more challenging for other uncommon viewpoints. Although uncommon viewpoints are less populated in most image datasets, cameras in embodied AI and robotics are often positioned in these uncommon viewpoints. Hence it is of crucial importance for LMMs to retain good 3D spatial reasoning performance for both common and uncommon viewpoints. To fairly compare the 3D spatial reasoning capabilities of LMMs w.r.t. different camera viewpoints, we annotate another 672 visual question-answer pairs on multi-view synthetic images rendered from the HSSD dataset [32].
@@
Besides benchmarking a wide variety of open-sourced and proprietary LMMs, our 3DSRBench serves as an important diagnosis benchmark for developing 3D spatially intelligent LMMs. Inspired by previous studies on 3D awareness of visual foundation models [20, 44], our 3DSRBench takes one step further and evaluates LMMs on fundamental 3D spatial reasoning questions, which provide valuable insights regarding the 3D awareness of visual encoders [47, 14, 25, 46, 34] and the 3D reasoning abilities of language models [18, 59, 53, 19]. Such results would shed light on downstream tasks that build on 3D spatial reasoning, such as automatic navigation and robotic manipulation.
@@
To enable a comprehensive and robust evaluation of 3D spatial reasoning abilities, 3DSRBench adopts several key designs: (1) balanced data distributions in multiple aspects, such as balanced answer distribution and complementary images pairs that lead to opposite answers given the same question (see Fig. 1(b)); (2) avoiding questions with shortcuts or trivial answers; and (3) a novel FlipEval strategy for robust evaluation of 3D spatial reasoning abilities.
@@
Our 3DSRBench significantly advances the evaluation of 3D spatial reasoning abilities and provide valuable findings and insights about the future development of LMMs. We benchmark a wide variety of open-sourced and proprietary LMMs on 3DSRBench and study their 3D spatial reasoning abilities w.r.t. different types of 3D awareness. We further investigate how various visual encoder designs and scaling of language model sizes can benefit 3D spatial reasoning abilities. Moreover, with the paired images in 3DSRBench-synthetic, we analyze the robustness of 3D spatial reasoning abilities w.r.t. uncommon camera 6D viewpoints. Lastly, by analyzing failure modes of state-of-the-art LMMs, we highlight limitations of current LMMs and discuss possible future improvements. Experimental results on different splits of our 3DSRBench provide valuable findings and insights that will benefit future research on 3D spatially intelligent LMMs.
@@ 2 Related Works
Spatial reasoning.
Early works [30, 28, 36, 31] studying spatial reasoning focused on spatial relationships w.r.t. the viewer, e.g. left/right relationships from the viewer’s perspective. We regard these as 2D spatial relationships as they can be derived merely from 2D bounding boxes of the objects. To study how LMMs can perceive and understand 3D spatial relationships, previous datasets often adopted synthetic environments, e.g., Blender, with controllable simulation and 3D groundtruths for automatic question-answer generation [15, 55, 56]. However, synthetic images in these datasets exhibit a large domain gap with natural images and it remains unclear if insights and findings from these datasets would generalize to the real image domain. More recent works, such as SpatialRGPT [17] and Cambrian-1 [52], built on existing datasets with 3D annotations [9, 50, 8, 22, 49, 11] and generated visual question-answer pairs with pre-defined rules. Despite the improved image quality, they are essentially limited to a small number of rigid object categories in Omni3D [9] and the automatically generated VQAs are subject to shortcuts and biases. To enable a comprehensive and robust evaluation of the 3D spatial reasoning capabilities, we manually annotate visual question-answer pairs on diverse and open-vocabulary entities, such as logos on a car or arrows on the billboard, enforcing balanced data distributions in multiple aspects and avoiding questions with shortcuts or trivial answers.
@@
In this section we introduce 3DSRBench for comprehensively analyzing the 3D spatial reasoning capabilities of LMMs. We start by presenting the design considerations in Sec. 3.1, i.e., how these design choices lead to a robust and valuable evaluation of 3D spatial reasoning capabilities. Then we show the four main question types in Sec. 3.2, as well as the challenges in each type of questions. Next we introduce the three splits of 3DSRBench and their scopes in Sec. 3.3. In Sec. 3.4 we present our evaluation strategies, including CircularEval and FlipEval. Please refer to Sec. A in supplementary materials where we provide details of our data collection and summary statistics of 3DSRBench.
@@ 3.1 Design of 3DSRBench
When developing 3DSRBench, we incorporate the following four key designs to enable a robust and valuable evaluation of 3D spatial reasoning capabilities. First, our 3D spatial reasoning questions are based on open-vocabulary entities. Previous spatial reasoning benchmarks [12, 52] largely relied on existing datasets with 3D annotations [9], which limited their scope to a small number of rigid object categories. In our 3DSRBench, we annotate 3D spatial reasoning questions across a broad range of open-vocabulary entities (see Fig. 1), enabling a thorough analysis of the 3D awareness and 3D reasoning capabilities of LMMs over diverse, commonly encountered real-world objects. Next, we avoid questions with shortcuts or trivial answers. For instance, objects higher in 3D space are usually higher in 2D space. We collect diverse VQAs and avoid those with clear shortcuts (see Fig. 2(a)). Also, when comparing which of the two objects has a smaller 3D distance to a third anchor object, we avoid the cases when there is a significant gap between the two distances, which lead to trivial answers. Moreover, we implement a balanced data distribution in various aspects, such as a roughly same number of yes/no answers and complementary image pairs [24] that lead to opposite answers given the same 3D spatial reasoning question (see Fig. 1(b)). This effectively removes priors in the answer distribution, e.g., pedestrians are often located lower than street lights, or the fact that objects higher in 3D space are also higher in 2D image plane. This design ensures that models cannot exploit biases or shortcuts for a higher benchmark performance. Lastly, we adopt special evaluation strategies for robust evaluation, including previous CircularEval [41] and our novel FlipEval (see Sec. 3.4).
@@ 3.2 Question Types
Multi-object reasoning questions consider the 3D spatial relationships between multiple objects, such as asking which side of an object is facing another object, or with three objects, asking which of the given objects is facing towards or closer to the third object. These questions require more advanced 3D awareness than simpler 3D concepts such as “closer” (to the camera) or “higher”, and require more complex 3D spatial reasoning, such as comparing distances between multiple objects from multi-step 3D computation.
(a) Height questions with different camera pitch rotations.
(b) Comparison between 2D and 3D spatial reasoning questions.
@@
Table 1: Overview of the 4 main types of 3D spatial reasoning questions and what kinds of 3D awareness and spatial reasoning are needed to answer each types of questions.
@@ 4 Experiments
We first introduce our experimental settings in Sec. 4.1. Next in Sec. 4.2 we benchmark various LMMs on 3DSRBench. We further study how various model designs, i.e., choice of visual encoders and scaling of language models, attribute to the 3D spatial reasoning abilities. Then we evaluate various LMMs on our 3DSRBench-synthetic and analyze the robustness of LMMs w.r.t. uncommon camera viewpoints in Sec. 4.3. Lastly we present some failure cases of GPT-4o and Gemini 2.0 in Sec. 4.4, highlighting limitations of current state-of-the-art LMMs and discussing possible future improvements.
@@ 4.1 Experimental Settings
With our 3DSRBench, we study: (1) standard 3D spatial reasoning abilities, by benchmarking various LMMs on 3DSRBench-real with VQAs on real images from MS-COCO [37], and (2) robustness of 3D spatial reasoning abilities w.r.t. uncommon camera viewpoints, by analyzing the performance gap between the two 3DSRBench-synthetic splits with common and uncommon viewpoints.
@@
We develop rule-based methods to augment the annotated visual question-answer pairs and obtain a larger number of testing data with a balanced and rich set of 3D spatial relationships. For instance, given a question asking which object is higher in the 3D world space, we generate a new question asking which object is lower in the 3D world space. We further adopt FlipEval that augments the question set by horizontally flipping the images. This leads to a total of 5,250 questions on MS-COCO images, i.e., 3DSRBench-real, and 1,692 questions on synthetic images, i.e., 3DSRBench-synthetic.
Evaluation.
To evaluate the correctness of free-form answers, we follow MMBench [41] and use exact matching to parse choice labels, or LLM-assisted evaluation, e.g., with gpt-4, when matching fails. We further adopt CircularEval [41] that repeats a question $N$ times, each with a different ordering of the choices. $N$ is the number of choices.
4.2 Results on 3D Spatial Reasoning Abilities